You build a box that is 4 feet long, 5 feet wide, and 8 feet high. The distance from the top left front corner to the top right back corner is ____ feet.
The distance from the bottom left front corner to the top right back corner is _____ feet.

Answer by mananth(16949)

the base is 4 feet by 5 feet.
the diagonal of the base has to be worked out.
using pythagoras theorem
4^2+5^2= diagonal ^
25+16=diagonal^2
sqrt39 = diagonal
..
the base diagonal, The distance from the bottom left front corner to the top right back corner, and the height form a right triangle.
sqrt((sqrt39)^2+8^2))= the required length
sqrt(39+64)
= sqrt (103)
=10.15 feet both will be same